OpenStudy (anonymous):

3x^2-6x=4 Pls help me find the solution of the quadratic equation by completing the square...

OpenStudy (anonymous):

Move all terms to one side 3x^2−6x−4=0

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt {b^2-4ac}}{2a} \] In this case, a=3, b=−6 and c=−4

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

Plug in the numbers and solve for x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[x=\frac{-(-6) \pm \sqrt {-6^2-4(3)(-4)}}{2(3)}\]

OpenStudy (anonymous):

that's as far as I can go, the rest is on you.

OpenStudy (anonymous):

:(( pls..

OpenStudy (wolf1728):

I think I could help :-)

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

Simplify: \[x=\frac{6+2\sqrt{21}}{6},\frac{6-2\sqrt{21}}{6}\] Answer: \[x=1+\frac{\sqrt{21}}{3},1-\frac{\sqrt{21}}{3}\]

OpenStudy (anonymous):

?-?

OpenStudy (anonymous):

@OmaiGod do you like math? do you know how to use basic arithmetic operations? basic equation manipulations?

OpenStudy (anonymous):

no.. help me pls..

OpenStudy (anonymous):

... it will be very hard for you to learn if you don't have any interest in this subject... what grade are you in?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok i will try what i can do for you to learn some of the basic concepts of quadratic equations...

OpenStudy (anonymous):

do you know what do numbers and letters represent in a equation?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

ok... numbers are called constants... if you say 3, it is a number 3, no other values...

OpenStudy (anonymous):

0k...

OpenStudy (anonymous):

for letters, we are using them as variables... they can have any value of number as long as they satisfy the quality in the equations...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you know linear equations?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

quadratic equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok let's proceed to completing the square...

OpenStudy (wolf1728):

I posted my solution by private message instead of interrupting this discussion. :-)

OpenStudy (anonymous):

what @TwinYang gave you is the Quadratic Formula used to find the solution of any quadratic equation...

OpenStudy (anonymous):

Quadratic Formula was also derived from completing the square...

OpenStudy (anonymous):

let's use the general equation of \(ax^2+bx+c=0\) and as @wolf1728 replied, he gave you the solution already...

OpenStudy (wolf1728):

yes a worked out solution

OpenStudy (anonymous):

still interested @OmaiGod ?

OpenStudy (anonymous):

yes..

OpenStudy (wolf1728):

I'm wondering just how familiar OmaiGod is with all this. Do you know the difference between a quadratic and a linear equation? Do you know how to graph an equation?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

do you know the degree of the equation?

OpenStudy (anonymous):

n00

OpenStudy (anonymous):

ok that is a problem too...

OpenStudy (anonymous):

you can determine the degree of the equation by noting the biggest exponent of the variable (letter used such as x) in the given simple equation...

OpenStudy (anonymous):

for \(x + 3=0\), what is the degree of the equation?

OpenStudy (anonymous):

a variable without an exponent is automatically raised to 1...

OpenStudy (anonymous):

degree of the equation also gives you an idea of the graph of equation...

OpenStudy (anonymous):

a degree 1 equation is also known as linear equation and represented by a straight line...

OpenStudy (anonymous):

a degree 2 equation is also known as quadratic equation and represented by a conics such as circle, parabola, ellipse, and hyperbola...

OpenStudy (anonymous):

still want me to continue @OmaiGod ? i think you need a lot of time to absorb all of these... it cannot be done in a single hour or class session...

OpenStudy (precal):

ok we are so off tangent I can help you by using completing the square it really is not so bad or hard

OpenStudy (precal):

3x^2-6x=4 ok first we need to factor out 3 3(x^2-2x)=4 if the 3 was not there, this would be an easier problem to do

OpenStudy (precal):

3(x^2-2x+______) --- 3 (____) =4 better to do it this way

OpenStudy (precal):

historically let me give you a quick reason why we complete the square

OpenStudy (precal):

this is x^2 (btw I am using algebra tiles to illustrate this)

OpenStudy (precal):